CAT Chat Minutes

June 6, 1997

 

Issues that require action and/or follow-up
1. I was wondering if the emergency exit doors in the LOMs will be locked soon?

This is tied to the card-key system. These door will be alarmed when the card-keys are operational.

2. Last week we had a student staying at the Guest House. The restaurant was closed early. The student had no car and it was very difficult for him to get any food. Could we get some type of list or could it be possible for restaurants to be able to deliver food on-site?

We will look into this and advise you.

3. The LOM 433 parking lot again has a large trash receptacle in it. This receptacle obstructs the view of the entrance to the parking lot. Can something be done?

The FC has had the receptacle moved. The dumpster will be in that location until sometime in July.

4. In 435 and 432, the "dollar" change machines are not working. Who do we call?

The FC has already taken care of this by calling the vendor. We will also post a phone number of the person to contact for repairs.

5. The cleaning people have been spraying the counters in our wet labs. They have sprayed on our equipment. Do they have a schedule that we could have so we would know when they will be cleaning these areas and we can be prepared for them? When I asked them not to do this, they indicated that they are told what to do by their supervisors and that if this work is not done, they will get in trouble.

We will look into this and try to find out if they are working on a schedule, and if we can get that schedule from them.

6. Now that we are using an automated system to fill the Dewars, can we rethink the safety procedures for this activity? How do we fill small Dewars? What is the schedule for refilling the storage tanks?

Steve D. and Bruce S. will check into the procedure changes. To fill the storage tanks, a team comes around and fills them on an as-needed basis. Small dewars can be filled using the appropriate attachment provided.

7. What is the status of the Liquid Nitrogen Distribution System?

A committee of CAT members was formed to address this. The design has been prepared, and cost estimates have been obtained by APS. We are waiting for the committee recommendations on implementation.

8. I found the rear entrance from the APS side to the Guest House closed when I returned at 0400 from the Experiment Floor. The only way I could get in was to retrace my path and enter from the front door. Can this inconvenience be corrected?

We have instructed Marriot to install locks (card reader), which you can open with a room key (card key) at night time. This should solve the problem. They may also install such a lock (card reader) on the front door of the Guest House.

Information
The gateway has been worked on. Some additional testing will be done on Monday. It appears to be working fine.

Real-time feedback is on schedule. We have planned some tests for Tuesday. This will run between .02 and 20 Hz and will be concurrent with the slow orbit correction. On Monday, we will send you a message with the PVs.

User Operations will continue through Monday 6-9 at 0800. At that time, we will have 24 hrs of Machine Studies, followed on 6-10 at 0800 with one shift of Shielding Verification, then User Operations will begin at 1600 and continue until 6-15.

We have delivered 81 hours of beam without a fault.

Secretary of Energy Frederico Pena will visit the APS on Wednesday. At this time, he is scheduled to be here ~ 9:00 a.m to 10:15 a.m. We will bring the Storage Ring down for his visit. He will make two stops in the Experiment Hall floor at Sectors 4 and 16. He will exit from 435. You may be requested to help with some of the housekeeping. We will advise you on Monday of the exact time the Storage Ring will be brought down. We expect the Storage Ring back in operation by 1100.

The Mutual CAT Safety Oversight Group will begin the review of SRI CAT on Monday, June 9, 1997. The CATs involved in this group are: BESSRC, CMC, DND and SRI. Chuck Wiley is chairing this group.

We have had some difficulties with beam stability in Sector 17, which have been corrected, and we are looking at similar problems in Sector 12, which appear to be in the SR.